# QUITE: A Query Rewrite System Beyond Rules with LLM Agents

### v1

@@ Abstract.
Query rewrite transforms SQL queries into semantically equivalent forms that run more efficiently. Existing approaches mainly rely on predefined rewrite rules, but they handle a limited subset of queries and can cause performance regressions. This limitation stems from three challenges of rule-based query rewrite: (1) it is hard to discover and verify new rules, (2) fixed rewrite rules do not generalize to new query patterns, and (3) some rewrite techniques cannot be expressed as fixed rules. Motivated by the fact that human experts exhibit significantly better rewrite ability but suffer from scalability, and Large Language Models (LLMs) have demonstrated nearly human-level semantic and reasoning abilities, we propose a new approach of using LLMs to rewrite SQL queries beyond rules. Due to the hallucination problems in LLMs, directly applying LLMs often leads to nonequivalent and suboptimal queries. To address this issue, we propose QUITE (query rewrite), a training-free and feedback-aware system based on LLM agents that rewrites SQL queries into semantically equivalent forms with significantly better performance, covering a broader range of query patterns and rewrite strategies compared to rule-based methods. Firstly, we design a multi-agent framework controlled by a finite state machine (FSM) to equip LLMs with the ability to use external tools and enhance the rewrite process with real-time database feedback. Secondly, we develop a rewrite middleware to enhance the ability of LLMs to generate optimized query equivalents. Finally, we employ a novel hint injection technique to improve execution plans for rewritten queries. Extensive experiments show that QUITE reduces query execution time by up to 35.8 % over state-of-the-art approaches and produces 24.1 % more rewrites than prior methods, covering query cases that earlier systems did not handle.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@ 1. Introduction
Inefficient SQL queries remain a long-standing challenge in many database applications. These inefficiencies often arise from poorly written queries, which are typically produced by inexperienced users or automatically generated by systems (Li, 2019; Zhou et al., 2020; Pirahesh et al., 1992). Query rewrite is a promising approach that transforms an original SQL query into a semantically equivalent form with significantly better performance (Finance and Gardarin, 1991a; Pirahesh et al., 1992; Wang et al., 2022). An effective rewrite can improve the execution time of a poorly written SQL query by several orders of magnitude (Wang et al., 2022; Zhou et al., 2021).
@@
Figure 1. An Example of Using Strategies beyond Rewrite Rules and Using Context-Aware Analysis. (b) boosts speed 37.5x over (a) by using a common table expression (CTE) to eliminate redundant calculations. (c) leverages data characteristics by joining sale (10 M rows) with prod (100 K rows) yields a 100 K-row intermediate (vs. 1.1 M rows in (b) ), achieving a 2.79× speedup.
Figure 2. An Example of Using Query Intent
However, even state-of-the-art query rewrite approaches can handle only a small subset of queries and may lead to performance regressions (Zhou et al., 2021; Li et al., 2024; Sun et al., 2024). This limitation stems from the intrinsic challenges of rule-based query rewrite. First, the rich features of SQL and the nuances of semantics make it rather difficult to discover and verify new rewrite rules (Pirahesh et al., 1992; Chu et al., 2017b; Wang et al., 2022). Even state-of-the-art rule discovery system (Wang et al., 2022) can only support a narrow subset of operators. Second, fixed rewrite rules rely on pattern matching and therefore are fundamentally unable to optimize unseen or complex query patterns (Leis et al., 2015; Bai et al., 2023; Dong et al., 2023). Third, many effective rewrite strategies cannot be expressed by rewrite rules (Dong et al., 2023). For example, the Common Table Expressions (CTE) conversion in Example 1.1 cannot be captured by rewrite rules. This is because such conversion requires explicit control over evaluation order and sub-plan sharing, which is beyond the scope of relational-algebra rewrite rules (Graefe, 1993; Wang et al., 2022; Begoli et al., 2018).
Given these limitations, we revisit the query rewrite problem from the perspective of human experts and identify three key capabilities that humans exhibit, but rule-based systems lack. (1) Humans can perform rewrites that follow rules but go beyond the expressiveness of pattern matching (Example 1.1). (2) Experts can apply intent-oriented rewrites that capture the underlying goal of the query, which are not easily expressed as explicit rules (Example 1.2). (3) Humans can conduct context-aware rewrites by considering the underlying data distributions and cost estimates provided by the query optimizer (Example 1.3).
@@
Query Rewrite based on Query Intent. Figure 2 (a) shows a real-world query from Apache Calcite (apa, 2021). Rule-based rewrite cannot optimize this query (Figure 2 (b). Identifying that the query intent is to sum salaries directly from the EMP table, human can simplify the original query by eliminating unnecessary joins and sub-queries that do not contribute to the intent of the query, achieving a 2.17 $\times$ execution improvement in our experimental setup.
@@
Context-Aware Query Rewrite. In figure 1 (c), instead of directly joining two large tables (the sale with 10 M rows and the ttime with 1.4 M rows), the query first joins the sale table with the prod table (100 K rows). This highly selective join would produce an intermediate result of only  100 K rows. Next we join this with ttime table and this yields a 2.79 $\times$ speedup over (b) in our experimental setup. Such rewrites, which reduce intermediate result size, are rare in rule-based systems and require analyzing data distribution and selectivity.
@@
To address the limitations of rule-based query rewrite, the poor scalability of manual rewrite, and the unreliable nature of LLMs, we propose QUITE, a training-free and feedback-aware system that smartly leverages LLMs to rewrite SQL queries into semantically equivalent forms with significantly improved performance. QUITE  supports a broader range of query patterns and rewrite strategies, while maintaining good scalability. First, to enable LLMs to capture query data characteristics and leverage database feedback, we design a multi-agent framework controlled by a finite state machine (FSM) to iteratively refine the rewrite process. We decompose the complex rewrite process into a set of subtasks, each handled by a specialized LLM agent to reduce hallucinations and improve rewrite ability. Second, to enhance the ability of LLMs to generate high-quality query rewrites, we provide a rewrite middleware that includes: (1) a structured knowledge base extracted from official documentation and query rewrite discussions on web forums, (2) a hybrid SQL corrector to ensure query equivalence, and (3) an agent memory buffer that manages important rewrite contexts. Third, we employ a novel hint injection technique to generate better execution plans for rewritten queries. This ensures that query execution exactly follows the optimized plan recommended by LLM agents, avoiding unintended modifications by the query optimizer due to inaccurate cost estimations. Additionally, this technique enables further improvements by incorporating physical-level optimizations.
@@ 2.1. Query Rewrite
Query rewrite is a process of transforming a query into another query that is semantically equivalent but with improved performance. It is a distinct step prior to query optimization, operating at the level of the user-accessible application programming interface (API) (e.g., SQL).
@@
Conventional studies treat query rewriting as a transformation between plain SQL statements, typically using rule-based methods. There are two research directions:
(1) Discovery of new rules. Approaches such as WeTune (Wang et al., 2022) automatically generate and verify logical plan transformations, but they are limited to certain operator types or algebraic equivalences. Techniques based on equality saturation (Nandi et al., 2021) and user-guided tools like QueryBooster (Bai et al., 2023) lack full autonomy and broad applicability. More recently, GENRewrite (Liu and Mozafari, 2024) uses LLMs to suggest rules in natural language. Yet GENRewrite must overcome challenges in formally verifying equivalence, preventing incorrect suggestions, and choosing the best candidates.
(2) Effective use of existing rules. A rule-based rewrite framework must answer two key questions: which rules to include in the rule set and how to select among applicable rules during rewriting. LearnedRewrite (Zhou et al., 2021) uses Monte Carlo Tree Search (MCTS) with learned cost models to explore the space of rule applications, but its performance depends on the accuracy of cost model estimation. LLM-R2 (Li et al., 2024) leverages LLMs for rule selection based on similar query examples. However, assembling its curated demonstration pool entails costly execution of rule combinations per query, and disparate data distributions between demonstration and target queries may lead to suboptimal results.
@@ 2.2. Query Hints
Query hints are SQL extensions that provide instructions to the database’s query engine to influence execution plan selection (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021). The following example uses the pg_hint_plan extension (NTT OSS Center DBMS Development and Support Team, 2012) to demonstrate how query hints operate:
@@
As shown above, query hints will provide precise control over the optimizer’s choice of physical operators and access methods: /*+ HashJoin(employees departments) */ compels a hash join for the $employees\bowtie departments$ operation, and /*+ SeqScan(employees)*/ enforces a full table scan of the employees relation, thereby overriding the optimizer’s default cost-based plan selection.
@@ 2.3. LLM and LLM Agents
LLM for Database Tasks. LLMs are increasingly applied to database tasks like text-to-SQL (Gao et al., 2023), SQL workload generation (Lao and Trummer, 2025), data processing (Albalak et al., 2024), and database optimization (Lao et al., 2024a; Zhao et al., 2025; Lao et al., 2024b; Tan et al., 2025; Lao et al., 2025), with recent work focusing on query optimization (Li et al., 2024; Sun et al., 2024; Liu and Mozafari, 2024). Although these methods show LLMs can understand and modify SQL, the inherent complexity of query rewrite presents substantial challenges.
LLM Agents. The primary advantage of LLM agents over LLMs is their ability to collect feedback from the execution environment (e.g., database execution metrics) and dadapt their subsequent reasoning and outputs accordingly. (Wu et al., 2025). They have proven particularly effective in complex problem solving—for example, SWE‐agent in software engineering (Yang et al., 2024) and ChemCrow in chemistry (Bran et al., 2023). QUITE applies this LLM agents methodology to query rewrite by gathering DBMS feedback and adjusting each query candidate accordingly. This performs a feedback-aware refinement, enabling reliable and advanced query rewrite that extends beyond predefined rules or the capabilities of standalone LLMs.
@@ 3. System Overview
In this section, we present our system’s end-to-end pipeline (Section 3.1) and describe its core components (Section 3.2).
3.1. System Pipeline
The system pipeline is illustrated in Figure 3. ❶ User provides the SQL query to be optimized along with the corresponding database configuration parameters (e.g., DBMS type, host address). ❷ QUITE obtains metadata (e.g., database statistics, query information) to generate a reasoning chain. This chain, which contains rewrite proposals and corresponding rewritten SQL candidates, is then used to generate the rewritten SQL for downstream enhancement and validation. ❸ QUITE consolidates the SQL candidates generated by the reasoning chains, constructs the prompt and selects and then applies global rewritten SQL enhancements. ❹ QUITE validates the syntactical correctness and semantic equivalence of the rewritten SQL. ❺ QUITE reassesses the rationality of the rewritten SQL to decide whether to proceed or to rewrite the SQL again referring to the rewrite history. In Query Hint Recommender, ❻ QUITE first reduces the hint search space by selecting and creating high-quality hints. Then, ❼ QUITE recommends and injects the most suitable query hints for the rewritten SQL. ❽ Finally, QUITE outputs the final rewritten SQL together with optimization proposals.
3.2. System Components
QUITE has three components: Rewrite Middleware, LLM Agent-Based Query Rewriter, and Query Hint Recommender. They work as follows:
LLM Agent-Based Query Rewriter. The rewriter utilizes a training-free and feedback-aware LLM agent-based Finite State Machine to rewrite queries in multiple stages. First, we formalize query rewrite as a Markov Decision Process (MDP), treating each SQL form as a state, rewrite operations as actions, and cost‐reduction estimates as rewards to guide the LLM’s chain‐of‐thought toward lower‐cost rewrites (see Subsection 5.1). Second, we decompose the rewrite workflow into an FSM with three specialized agents, utilizing a multi‐agent collaboration strategy and constructing an FSM to ensure a stable workflow (see Subsection 5.2).
Rewrite Middleware. The toolkit prepares three specialized tools to enhance the rewrite capability of LLM agents. First, we construct a structured knowledge base using a multi-source query rewrite knowledge integration pipeline to guide the LLM agent to generate high-quality rewrite proposals (see Section 5.1). Second, we develop a hybrid SQL corrector incorporating an LLM agent to validate and correct the syntactical correctness and semantic equivalence of the rewritten SQL (see Section 5.2). Third, we design an agent memory buffer that can extract key information from previous agent conversations, thus reducing agent communication overhead and LLM hallucination in the rewrite process (see Section 5.3).
Query Hint Recommender. The recommender explores the query hints in two steps. First, we carefully construct a promising query hint base, thereby reducing the search space (see Section 6.1). Second, we select appropriate hints for the rewritten query to improve its execution plan (see Section 6.2). Finally, the selected hints are injected into the rewritten SQL to produce the final output.

### v2

@@ Abstract.
Query rewrite transforms a given SQL query into a semantically equivalent query that can be executed more efficiently. Existing approaches mainly rely on predefined rewrite rules. However, they can only handle a small subset of queries and may lead to performance regressions. This limitation arises from the intrinsic challenges of rule-based query rewrite: (1) it is hard to discover and verify new rewrite rules, (2) fixed rewrite rules do not generalize to new query patterns and are insufficient for handling complex queries, and (3) some rewrite techniques cannot be expressed by fixed rules. Motivated by the fact that human experts exhibit significantly better rewrite ability but suffer from scalability, and Large Language Models (LLMs) have demonstrated nearly human-level semantic and reasoning abilities, we propose a new approach of using LLMs to rewrite SQL queries beyond rules. Due to the hallucination problems in LLMs, directly applying LLMs often leads to nonequivalent and suboptimal queries. To address this issue, we propose QUITE (query rewrite), a training-free and feedback-aware system based on LLM agents that rewrites SQL queries into semantically equivalent forms with significantly better performance, covering a broader range of query patterns and rewrite strategies compared to rule-based methods. Firstly, we design a multi-agent framework controlled by a finite state machine (FSM) to equip LLMs with the ability to use external tools and enhance the rewrite process with real-time database feedback. Secondly, we develop a rewrite middleware to enhance the ability of LLMs to generate optimized query equivalents. It includes a structured knowledge base for domain knowledge, an SQL corrector to ensure equivalence, and an agent memory buffer to manage critical rewrite contexts. Finally, we employ a novel hint injection technique to produce better execution plans for rewritten queries. Extensive experiments show that our method achieves up to a 35.8% reduction in query execution over state-of-the-art approaches. Moreover, it delivers 24.1% additional rewrites compared to prior methods, which cover query cases that earlier approaches could not handle.
@@ 1. Introduction
Inefficient SQL queries remain a long-standing challenge in many database applications, often arising from poorly written queries by inexperienced users or automatically generated systems (Li, 2019; Zhou et al., 2020; Pirahesh et al., 1992). Query rewrite is a promising approach that transforms an original SQL into a semantically equivalent form with better performance (Finance and Gardarin, 1991a; Pirahesh et al., 1992; Wang et al., 2022). An effective rewritten query can improve the execution time of a poorly written SQL query by several orders of magnitude (Wang et al., 2022; Zhou et al., 2021).
@@
Figure 1. An Example of Using Strategies Beyond Rewrite Rules (Example 1.1) and Using Context-Aware Analysis (Example 1.2)
Figure 2. An Example of Using Query Intent (Example 1.3)
However, the above approaches can handle only a small subset of queries and may lead to performance regressions (Zhou et al., 2021; Li et al., 2024; Sun et al., 2024). This limitation stems from the intrinsic challenges of rule-based query rewrite. First, the rich features of SQL and the nuances of semantics make it rather difficult to discover and verify new rewrite rules (Pirahesh et al., 1992; Chu et al., 2017b; Wang et al., 2022). Even state-of-the-art rule discovery system (Wang et al., 2022) can only support a narrow subset of operators. Second, fixed rewrite rules rely on pattern matching and therefore are fundamentally unable to optimize unseen or complex query patterns (Leis et al., 2015; Bai et al., 2023; Dong et al., 2023). Third, many effective rewrite strategies cannot be expressed by rewrite rules (Dong et al., 2023). For example, the Common Table Expressions (CTE) conversion in Example 1.1 cannot be captured by rewrite rules. This is because such conversion requires explicit control over evaluation order and sub-plan sharing, which is beyond the scope of relational-algebra rewrite rules (Graefe, 1993; Wang et al., 2022; Begoli et al., 2018).
Given these limitations, we revisit the query rewrite problem from the perspective of human experts and identify three key capabilities that humans exhibit, but rule-based systems lack. (1) Humans can perform rewrites that follow rules but go beyond the expressiveness of pattern matching (Example 1.1). (2) Humans can conduct context-aware rewrites by considering the underlying data distributions and cost estimates provided by the query optimizer (Example 1.2). (3) Experts can apply intent-oriented rewrites that capture the underlying goal of the query, which are not easily expressed as explicit rules (Example 1.3).
@@
Context-Aware Query Rewrite. In Figure 1 (c), instead of directly joining two large tables (the sale with 10 M rows and the ttime with 1.4 M rows), we first joins the sale table with the prod table (100 K rows). This highly selective join would produce an intermediate result of only  100 K rows. Next we join this with ttime table and yield a 2.79 $\times$ speedup over (b) in our experimental setup. Such rewrites, which reduce intermediate size, are rare in rule-based systems and require analyzing data distribution and selectivity.
@@
Query Rewrite based on Query Intent. Figure 2 (a) shows a real-world query from Apache Calcite (apa, 2021). Rule-based rewrite cannot optimize this query (Figure 2 (b)). Identifying that the query intent is to sum salaries directly from the EMP table, human can simplify the original query by eliminating unnecessary joins and sub-queries that do not contribute to the intent of the query, achieving a 2.17 $\times$ execution improvement in our experimental setup.
@@
To address the limitations of rule-based query rewrite, the poor scalability of manual rewrite, and the unreliable nature of LLMs, we propose QUITE, a training-free and feedback-aware system that smartly leverages LLMs to rewrite SQL queries into semantically equivalent forms with significantly improved performance. QUITE supports a broader range of query patterns and rewrite strategies, while maintaining good scalability. First, to enable LLMs to capture query data characteristics and leverage database feedback, we design a multi-agent framework controlled by a finite state machine (FSM) to iteratively refine the rewrite process. We decompose the complex rewrite process into a set of subtasks, each handled by a specialized LLM agent to reduce hallucinations and improve rewrite ability (addressing C1, C2). Second, to enhance the ability of LLMs to generate high-quality query rewrites, we provide a rewrite middleware that includes: (1) a structured knowledge base extracted from official documentation and query rewrite discussions on web forums, (2) a hybrid SQL corrector to ensure query equivalence, and (3) an agent memory buffer that manages important rewrite contexts (addressing C1, C2). Third, we employ a novel hint injection technique to generate better execution plans for rewritten queries. This ensures that query execution exactly follows the optimized plan recommended by LLM agents, avoiding unintended modifications by the query optimizer due to inaccurate cost estimations. Additionally, this technique enables further improvements by incorporating physical-level optimizations (addressing C2).
@@ 2.1. Query Rewrite
Query rewrite is a process of transforming a query into another query that is semantically equivalent but with improved performance. It is a preliminary step prior to query optimization, operating at the level of the user-accessible application programming interface (API) (e.g., SQL).
@@
Conventional studies treat query rewrite as a transformation between plain SQL statements, typically using rule-based methods. There are two lines of work:
(1) Discovery of new rules. Approaches such as WeTune (Wang et al., 2022) automatically generate and verify logical plan transformations, but they are limited to specific operator types and algebraic equivalences. Equality saturation (Nandi et al., 2021) and user-guided tools like QueryBooster (Bai et al., 2023) lack full autonomy and broad applicability. Recently, GenRewrite (Liu and Mozafari, 2024) uses LLMs to suggest rules in natural language. However, it still struggles with optimal rewrite selection because executing all candidates to measure latency incurs prohibitive overhead and its off-the-shelf verifiers permit equivalence checking for only a subset of rewrite pairs (Liu and Mozafari, 2024; Dong et al., 2023).
(2) Effective use of existing rules. A rule-based rewrite framework must answer two key questions: which rules to include in the rule set and how to select among applicable rules during rewriting. LearnedRewrite (Zhou et al., 2021) uses Monte Carlo Tree Search (MCTS) with learned cost models to explore the space of rule applications, but its performance depends on the accuracy of cost model estimation. LLM-R2 (Li et al., 2024) uses a trained selection model to retrieve the past rewrite query and rule demonstrations similar to the input SQL, and leverages them as an in-context prompt for the LLM to select promising rewrite rules. However, it still incurs high costs by executing every query demonstration on each rule set and faces distribution shifts between the demonstration and input SQL, since structurally similar queries may differ in row count, leading to suboptimal rule selection.
@@ 2.2. Query Hints
Query hints are SQL extensions that provide instructions to the database’s query engine to influence the selection of execution plans (Marcus et al., 2021; Anneser et al., 2023; Woltmann et al., 2023; Negi et al., 2021). These hints allow users to exert precise control over how the database optimizer chooses physical operators and access methods.
As shown below for pg_hint_plan extension (NTT OSS Center DBMS Development and Support Team, 2012), the hint /*+ HashJoin(employees departments) */ forces the query engine to use a hash join for the $employees\bowtie departments$ operation, while the /*+ SeqScan(employees)*/ enforces a full table scan of the employees relation. These hints can override the optimizer’s default cost-based plan selection.
@@ 2.3. LLM and LLM Agents
LLM for Database Tasks. LLMs are increasingly applied to database tasks like text-to-SQL (Gao et al., 2023), SQL workload generation (Lao and Trummer, 2025), data processing (Albalak et al., 2024), and database optimization (Lao et al., 2024a; Zhao et al., 2025; Lao et al., 2024b; Tan et al., 2025; Lao et al., 2025), with recent work focusing on query optimization (Li et al., 2024; Sun et al., 2024; Liu and Mozafari, 2024). Although these methods show that LLMs can understand and modify SQL, the inherent complexity of query rewrite presents substantial challenges.
LLM Agents. The primary advantage of LLM agents over LLMs is their ability to collect feedback from the execution environment (e.g., database execution metrics) and dadapt their subsequent reasoning and outputs accordingly (Wu et al., 2025). They have proven particularly effective in complex problem solving—for example, SWE‐agent in software engineering (Yang et al., 2024) and ChemCrow in chemistry (Bran et al., 2023). QUITE applies this LLM agents methodology to query rewrite by gathering DBMS feedback and adjusting each query candidate accordingly. This performs a feedback-driven refinement, enabling reliable and advanced query rewrite that extends beyond predefined rules or the capabilities of standalone LLMs.
2.4. Comparison with Other Works
Instead of focusing on finding optimal rewrite rules with limited coverage (Zhou et al., 2021; Li et al., 2024; Sun et al., 2024), QUITE addresses the core aspects of query rewriting by identifying semantic and physical rewrite bottlenecks through real-time interaction between LLM agents and the database, enhancing the detection of performance issues in the input SQL. Furthermore, QUITE generates a reasoning chain with natural language rewrite proposals and rewrites the SQL through multi-agent collaboration within a carefully designed Finite State Machine (FSM) pipeline and agent tools, ensuring equivalence and effectiveness. This approach enhances adaptability and efficiency in context-aware rewrites, compared to rule-based systems that miss complex query patterns and context-specific optimizations.
In contrast to GenRewrite (Liu and Mozafari, 2024), which directly utilizes LLMs to generate rewrite rules for rewriting, QUITE considers key factors such as the SQL’s physical plan, the database distribution, and the domain expert knowledge for analysis, ensuring better alignment with the execution environment. Additionally, QUITE introduces a specialized reasoning agent, which forms the core of its multi-agent framework, incorporating a planning mechanism to sequence high-impact rewrite actions, maximizing performance improvements and preventing inefficiencies and regressions.
Furthermore, QUITE incorporates a fine-grained hint injection technique into query rewrite process to further improve the query execution performance. To our knowledge, no other query rewrite systems have supported for these components.
@@ 3. System Overview
We present the system architecture of QUITE  as shown in Figure 3. QUITE comprises three components: LLM Agent-Based Query Rewriter, Rewrite Middleware, and Query Hint Recommender.
LLM Agent-Based Query Rewriter. The rewriter uses a training-free and feedback-aware LLM agent-based FSM to rewrite queries in multiple stages. First, ❶ a user provides the input SQL along with the corresponding database configuration (e.g., DBMS type, host address). Second, ❷ The rewriter obtains metadata (e.g., database statistics, query information) to generate a reasoning chain, which includes rewrite proposals and rewritten SQL candidates. This chain is then used to generate the rewritten SQL for further enhancement and validation. Next, ❸ we consolidate the SQL candidates to construct the LLM prompt, and apply global SQL enhancements to find the optimal rewritten SQL. Finally, ❹ we validate the syntactical correctness and semantic equivalence of the rewritten SQL, and ❺ reassess the rationality of the rewritten SQL to decide whether to proceed or rewrite the SQL again based on the rewrite history.
Rewrite Middleware. The toolkit provides three specialized tools to boost LLM rewrite ability. First, we construct a structured knowledge base using a multi-source integration to guide high-quality rewrite proposals (see Section 5.1). Second, a hybrid SQL corrector then validates the syntactical correctness and semantic equivalence of each candidate rewrite, reassessing rationality against rewrite history and triggering further refinement when necessary (see Section 5.2). Third, we design an agent memory buffer that captures key context from previous interactions to reduce communication overhead and LLM hallucination (see Section 5.3).
Query Hint Recommender. The recommender explores the query hints in two steps. ❻ First, we carefully construct a promising query hint base, thereby reducing the hint search space and enabling faster, more accurate hint selection (see Section 6.1). ❼ Second, we leverage LLM agent with data statistics to select appropriate fine-grained hints for the rewritten query to improve its execution performance (see Section 6.2). ❽ Finally, the selected hints are injected into the rewritten SQL to produce the final output.